3.Lockard Company purchased machinery on January 1, 2010 for 80,000. The machinery is estimated to have a salvage value of $8,000 after a useful life of 8 years. A) Compute 2010 depreciation expense using the straight-line method. B) Compute 2010 depreciation using method assuming the machinery was purchased on September 1,2010.



4.Use the information for Lockard Company given in 3. A) Compute 2010 depreciation expense using the sum-of the years ‘-digits method. B) Compute 2010 depreciation expense using the sum-of-the-years’-digits method assuming the machinery was purchased on April 1,2010

5. Use the information for Lockard Company given in 3. A) compute 2010 depreciation expense using the double-declining-balance method. B) Compute 2010 depreciation expense using the double-declining-balance method assuming the machinery was purchased on October 1,2010.
Answer #1

A) Computing the depreciation expense using Straight-line method:

The formula for calculating straight line depreciation is:

SLD = (Cost of the asset - Salvage value) / Useful life of the asset
= ($80,000 - $8,000) / 8
= $9,000

Therefore, the 2010 depreciation expense using straight line method is $9,000

B) If the asset is used only for a prt of year, th anual depreciation if prorated. The depreciation for the year ending Dec 31st would be :

First-year partial depreciation = $9,000 * (4/12)
= $3,000

The deprecaition expense is calculated only for the last four months from Sep to Dec in the first year.

Therefore, the depreciation expense assuming the machinery was purchased on Sep 1st is $3,000.

4.
A) Computing the depreciation expense using sum-of-years digits method:

Sum of years of useful life = [N (N + 1)] / 2

where N = 8yrs

Sum of years of useful life = 8 (9) / 2
= 36

uploaded image

B) If the asset is purchased on April 1st, then the first year partial depreciation is calculated as:

First year partial depreciation = $16,000 * (9/12)

                                             = $12,000

Second year depreciation = (3/12 x 8/36 x $72,000) + (9/12 x 7/36 x $72,000)

                                      = $4,000 + $10,500

                                      = $14,500

5. A) The doouble declining depreciation expense is calculated as:

Double declining depreciation = 2(Straight-line depreciation)

                                            = 2 (100% / 8)

                                            = 2 (12.5)

                                           = 25%

 

uploaded image

In this method, the residual value is not considered. However, the asset should not be declined below its estimated residual value. That is why we have not used double declining method for the last year.

If the asset is purchased on Oct 1st,

First-year partial depreciation = $20,000 * (3/12) 

                                             = $5,000

Second-year partial depreciation = [25% ($80,000 - $5,000)]

                                                 = $18,750